# Bayesian Deep-stacking for High-energy Neutrino Searches

### v1

@@ 1 Introduction
It has been more than a decade since IceCube identified a flux of high-energy neutrinos of cosmic origin [1]. While the first neutrinos were originally detected as a quasi-diffuse flux without a known origin, since then multiple individual neutrino sources have been identified, including the blazar TXS 0506+056 [2, 3] and the active galaxy NGC 1068 [4]. There is a growing number of additional blazars and non-blazar active galactic nuclei that have associated with detected neutrinos (e.g., [5, 6, 7, 8]). Despite these identifications, the vast majority of observed astrophysical neutrinos have still no known. It also appears that new source types may be necessary to explain the entire observed cosmic neutrino flux [9].
Due to the weakness of individual neutrino sources, identifications so far required multi-messenger information. The first identification, TXS 0506+056, began with an energetic neutrino trigger and searched for possible sources that directionally and temporally overlapped with the neutrino. Some other searches were carried out over a limited number of candidate sources, such as the brightest gamma-ray sources (e.g., [4]). Further search sensitivity could be gained by using not just information of the source position, but also incorporating information on the expected neutrino flux, estimated from the candidates’ electromagnetic emission, as a prior (e.g., [7, 8]).
@@
Bayesian models have previously been proposed for neutrino analyses. [12] implemented a Bayesian hierarchical model to place constraints on astrophysical neutrino populations. To limit computational cost and utilize publicly available data, this model built on the reconstructed flux and spectral index for astrophysical sources. [13] proposed a more general approach that directly incorporated the properties of each detected neutrino in a hierarchical Bayesian model, and demonstrated the applicability of such an approach to a population of sources. This latter work targets a similar direction as the present paper, with similar likelihood functions (c.f. Eq. 2.3 and their Eq. 5). While [13] focuses on the practical implementation of a Bayesian stacking method that is then tested on several simple examples, the present paper focuses more on (i) the conceptual introduction of Bayesian stacking, (ii) on giving a prescription on addressing questions relevant for deep-stacking analyses, and on (iii) carrying out a full-scale simulation to demonstrate the power of Bayesian inference on a realistic problem in comparison to frequentist methods.
@@ 2.1 Types of information
Generally, in a stacking analysis, there are four types of information. First, we have a set of observations, in our particular case a set of $n_{\nu}$ detected neutrino events. We will denote all reconstructed properties of neutrino $i$ with $x_{i}$. These can include its reconstructed energy $E_{i}$, reconstructed sky location $\omega_{i}$ and directional uncertainty $\sigma_{i}$, i.e. $x_{i}=\{{E}_{i},{\omega}_{i},\sigma_{i}\}$. Other, or different properties are also possible, for example other proxies for neutrino energy. The ensemble of observed neutrino data in the analysis will be denoted with $\mathbf{{x}}\equiv\{{x}_{1},{x}_{2},...,{x}_{N_{\nu}}\}$.
The remaining three types of information are parameters independent of neutrino observations, in our case source properties, which we will collectively denote with $\boldsymbol{\xi}$. Within this category, our second type of information is a set of source properties $\xi_{\rm fix}$ that are assumed to be fixed (known). These can include the sources’ sky location, luminosity distance, etc. These do not need to be specific values, but can be described by probability densities as well. The point is that we do not aim to learn about these values from the neutrino data.
Third, we have a set of source properties $\xi_{\rm nui}$ that are nuisance parameters. We do not know the values of these properties and we estimate them based on the observed neutrino data, but these are ultimately not of interest to us, and therefore we marginalize over them. For example, if our specific goal with an analysis is to establish the presence of a neutrino source population, we want to marginalize over all unknown source parameters in order to establish an overall detection likelihood.
Finally, the fourth category is a set of source properties $\xi_{\rm var}$ which are unknown and which we intend to uncover with the analysis. For example, we may be interested in the average luminosity of a neutrino source population, in which case we will keep the average luminosity as the variable of interest.
In the categorization of the properties that are either of interest or are marginalized over, it is typically beneficial to consider group-level parameters, or so-called hyperparameters, that influence the individual properties of all our sources. The average source luminosity is one example of such a hyperparameter. The use of hyperparameters has multiple advantages: (i) since there is more observational data affected by a hyperparameter than for a parameter of an individual source, we are typically better able to constrain it; (ii) since we can describe the key features of a source population with a relatively small number of hyperparameters as opposed to the parameters describing all individual sources, we can reduce the chance of overfitting; (iii) hyperparameters allow for the population level modeling of the sources. Bayesian inference that includes hyperparameters is referred to as hierarchical Bayesian inference, which we focus on in the following.
@@ 2.6 Detection through model comparison
A common method to interpret the odds ratio is to consider threshold values over which it is considered significant preference for one of the other model. A downside of this interpretation, however, is that such an interpretation cannot be directly compared to common frequentist significance estimates, such as the false alarm rate. In cases in which such a comparison is beneficial, one can consider the odds ratio as a test statistic and compute its distribution for a background-only case as comparison.
@@ 3.2 Average luminosity
Figure 1: Reconstructed average luminosity $\langle L_{0}\rangle$ as a function of the expected number of detected astrophysical neutrinos per source. The simulation used 1000 cataloged sources and a background of $1$ detected neutrinos per source on average.
@@ 4.3 Frequentist likelihood analysis with source normalization
The likelihood is maximized with respect to the $N_{\rm s}$ parameters $\hat{n}_{j}^{*}$, which can become quite numerous in deep-stacking scenarios, and become challenging computationally.
@@ 4.4 Bayesian analysis
$\pi(\hat{n}_{j}^{*})=\frac{1}{\sqrt{2\pi}\hat{n}_{j}^{*}\sigma_{S}}\exp\left(-\frac{(\log_{10}\hat{n}_{j}^{*}-\log_{10}\hat{n}_{j})^{2}}{2\sigma_{S}^{2}}\right)\;\;,\mathrm{and}\;\;\pi(\boldsymbol{0})=0$ (4.5)
@@
For large numbers of sources, solving the $N_{S}$-dimensional integral in the odds ratio is generally computationally expensive. However, in the use case considered here, the signal PDF in the likelihood drops quickly with increasing distance from the source. Each event will only contribute to the signal PDF of a few sources, and the signal PDF of sources that are not close to the event direction will be negligible. This allowed us to reduce the dimensionality of the integral by only considering the source that is the closest to the event direction. This is justified in the case that the average distance between sources is much larger than the assumed angular resolution of our hypothetical detector and reduces the calculation to the odds ratio to solving a series of one-dimensional integrals. Using more advanced integration techniques, such as Monte Carlo integration, the odds ratio can be calculated efficiently even for a large number of sources.
@@ 4.5 Results
Figure 2: TS distributions for different types of stacking analyses performed on simulations of the reference population in Section 4. Each row shows a different scenario of scatter between the neutrino luminosity injected in the pseudo-data and the one assumed in the likelihood analysis (no scatter, 0.3 dex, 0.5 dex and 1.0 dex). The columns show the three analysis approaches tested here: standard, source normalization, and Bayesian deep-stacking. In each panel, the TS distribution of the null hypothesis (blue, background only) is compared to the median of the TS distribution for the signal hypothesis (orange, background and sources). The shown p-value indicates the level of significance achieved for rejecting the null hypothesis on 50% of the pseudo-data samples (TSmed, red dashed line). Except where the distributions overlap sufficiently, they are obtained by extrapolating the background-only TS distribution with an appropriate fuction (green line, see text for details).

### v2

@@ 1 Introduction
It has been more than a decade since IceCube identified a flux of high-energy neutrinos of cosmic origin [1]. While the first neutrinos were detected as a quasi-diffuse flux without known origin, since then multiple individual neutrino sources have been identified, including the blazar TXS 0506+056 [2, 3] and the active galaxy NGC 1068 [4]. A growing number of additional blazars and non-blazar active galactic nuclei have also been associated with detected neutrinos (e.g., [5, 6, 7, 8]). Despite these identifications, the vast majority of astrophysical neutrinos still have no known origin, and additional source types may be necessary to explain the entire observed cosmic neutrino flux [9].
Due to the weakness of individual neutrino sources, previous identifications have typically required multi-messenger information. The first identified source, TXS 0506+056, was found by searching for electromagnetic counterparts temporally and directionally coincident with an energetic neutrino. Other searches considered limited candidate samples, such as bright gamma-ray emitters (e.g., [4]). Further sensitivity improvements were achieved by incorporating prior expectations of neutrino flux based on electromagnetic emission (e.g., [7, 8]).
Further search sensitivity could be gained by using not just information of the source position, but also incorporating information on the expected neutrino flux, estimated from the candidates’ electromagnetic emission, as a prior (e.g., [7, 8]).
@@
Bayesian models have previously been used in astrophysical analyses, for example in gravitational wave population inference (e.g., [12, 13, 14, 15]). Similar models have also been previously proposed for neutrino analyses. [16] implemented a Bayesian hierarchical model to place constraints on astrophysical neutrino populations. To limit computational cost and utilize publicly available data, this model built on the reconstructed flux and spectral index for astrophysical sources. [17] proposed a more general approach that directly incorporated the properties of each detected neutrino in a hierarchical Bayesian model, and demonstrated the applicability of such an approach to a population of sources. This latter work targets a similar direction as the present paper, with similar likelihood functions (c.f. Eq. 2.3 and their Eq. 5).
The aim of this paper is to facilitate the use of Bayesian deep-stacking methods in neutrino multi-messenger searches and parameter estimation. To achieve this, we first provide a clear conceptual introduction to Bayesian stacking, along with a detailed prescription for addressing key questions specific to deep-stacking analyses. We then perform a comprehensive simulation to demonstrate the advantages of Bayesian inference in a realistic scenario, explicitly comparing its performance to traditional frequentist approaches. Compared to previous hierarchical Bayesian analyses in neutrino astrophysics [16, 17], our primary novel contributions include explicitly addressing uncertainties in individual source luminosities and showcasing hierarchical Bayesian parameter estimation in a realistic context, directly contrasted with standard frequentist methods.
@@ 2.1 Types of information
We have the following types of parameters.
Neutrino Observation Parameters ($\mathbf{x}$).
These parameters are directly obtained from neutrino detection data. For detected neutrino event $i$, we denote these parameters as $x_{i}$, typically including the reconstructed neutrino energy $E_{i}$, the reconstructed sky location $\omega_{i}$, and the directional uncertainty $\sigma_{i}$. The neutrino observational data ensemble is then given by $\mathbf{x}\equiv\{x_{1},x_{2},\dots,x_{N_{\nu}}\}$. Alternative neutrino event properties or other proxies (e.g., energy estimators) can similarly be included depending on dataset specifics.
Fixed Source Parameters ($\xi_{\rm fix}$).
These are source properties assumed to be known or constrained independently from non-neutrino observations. Examples include source sky locations, luminosity distances, or host galaxy redshifts. These parameters are either assigned specific values or described by probability distributions derived from electromagnetic observations or theoretical models. Importantly, their values are not inferred from neutrino data and remain fixed throughout our inference.
Nuisance Parameters ($\xi_{\rm nui}$).
These represent unknown source properties that are not of primary interest in the analysis, but must be accounted for. Examples include individual source luminosities or intrinsic source spectral shape parameters if these are not the direct targets of the analysis. Since our primary analysis goals do not require the determination of these parameters, we marginalize over them, incorporating their uncertainties without explicitly reporting their inferred values.
Population (Hyper-)Parameters ($\xi_{\rm var}$).
These parameters characterize the neutrino source population rather than individual sources, and are explicitly targeted by the analysis. Examples include the average luminosity of the source population, intrinsic source rate densities, or parameters describing the population-level spectral distribution. The use of hyperparameters offers several advantages: (i) they influence multiple observations simultaneously, providing stronger statistical constraints; (ii) they reduce the complexity of the model, mitigating overfitting; and (iii) they yield insights into population-level properties rather than individual events.
Bayesian inference that includes hyperparameters is referred to as hierarchical Bayesian inference, which we focus on in the following.
@@ 2.6 Detection through model comparison
A common method to interpret the odds ratio is to consider threshold values over which there is a significant preference for one model over the other. While the odds ratio naturally lends itself to a Bayesian interpretation–namely, directly providing the probability that one model is favored given the observed data–such an interpretation cannot be directly compared to widely-used frequentist significance metrics, in particular the false alarm rate, that are standard in neutrino and multi-messenger astrophysics. To facilitate this comparison and ensure that our results align with established frequentist practices, we instead use the odds ratio as a test statistic and derive its distribution under the background-only scenario. This allows our results to be interpreted within a frequentist framework, thus enabling a direct comparison with existing frequentist significance estimates in the literature.
@@ 3.2 Average luminosity
Figure 1: Comparison of the reconstructed average source luminosity, $\langle L_{0}\rangle$, to the true luminosity, $L_{0}$, as a function of the expected number of detected astrophysical neutrinos per source. (top) reconstructed luminosity normalized by the true luminosity (the dashed line represents perfect reconstruction). (middle) absolute deviation of the reconstructed luminosity from the true value. (bottom) size of the 68% credible interval (CI) for the Bayesian method. The analysis assumes 1000 cataloged neutrino sources and an average background level of 1 neutrino event per source. Results from Bayesian inference (blue) and maximum likelihood estimation (orange) methods are shown for comparison.
@@ 4.3 Frequentist likelihood analysis with source normalization
The likelihood is maximized with respect to the $N_{\rm s}$ parameters $\hat{n}_{j}^{*}$, which can become quite numerous in deep-stacking scenarios, and become challenging computationally. To match this challenge, the likelihood calculations for this manuscript have been optimized using the PyTorch [^3] https://pytorch.org package that provides analytic gradients for the likelihood minimization and supports multi-core architectures as well as calculations of the likelihoods on specialized hardware such as GPUs.
@@ 4.4 Bayesian analysis
$\pi(\hat{n}_{j}^{*})=\frac{1}{\sqrt{2\pi}\hat{n}_{j}^{*}\sigma_{S}}\exp\left(-\frac{(\log_{10}\hat{n}_{j}^{*}-\log_{10}\hat{n}_{j})^{2}}{2\sigma_{S}^{2}}\right)\;\;,\mathrm{and}\;\;\pi(\boldsymbol{0})=1$ (4.5)
@@
For large numbers of sources, solving the $N_{S}$-dimensional integral in the odds ratio is generally computationally expensive. However, in the use case considered here, the signal PDF in the likelihood drops quickly with increasing distance from the source. Each event will only contribute to the signal PDF of a few sources, and the signal PDF of sources that are not close to the event direction will be negligible. This allowed us to reduce the dimensionality of the integral by only considering the source that is the closest to the event direction. This is justified in the case that the average distance between sources is much larger than the assumed angular resolution of our hypothetical detector and reduces the calculation of the odds ratio to calculating a series of one-dimensional integrals, which we perform using simple Gaussian quadrature (QUADPACK [20]).
Therefore our simulation code is currently limited to the conditions described above and does not scale favorably to even larger numbers of sources, when the average distance between sources approaches the angular resolution of the instrument. Multi-dimensional integrations are unavoidable in this case, and more advanced integration techniques, such as Monte Carlo integration, need to be evaluated and implemented to efficiently calculate the odds ratio in this case.
@@ 4.5 Results
Figure 2: Test statistic (TS) distributions for different stacking analysis approaches applied to simulated data from the reference neutrino source population described in Section 4. Rows correspond to different scenarios of scatter between the neutrino luminosity injected into the simulated data and the luminosity assumed in the likelihood analysis: no scatter, 0.3 dex, 0.5 dex, and 1.0 dex, from top to bottom respectively. Columns compare three analysis methodologies: maximum likelihood with sample normalization (left), maximum likelihood with source normalization (center), and Bayesian deep stacking (right). Each panel shows the TS distribution for the background-only (null) hypothesis in blue, and the median TS distribution under the signal-plus-background hypothesis in orange. The p-value indicated in each panel measures the statistical significance of rejecting the null hypothesis at the median TS of the signal distribution ($\mathrm{TS}_{\mathrm{med}}$, indicated by the red dashed line). When necessary, the significance is computed by extrapolating the background-only TS distribution using an appropriate fit (green line, see main text for details).